# Homogeneous $\mathrm{G}_{2}$ and Sasakian instantons on the Stiefel $7$-manifold

### v1

@@ Abstract
We study homogeneous instantons on the seven dimensional Stiefel manifold $V^{5,2}$ in the context of $\mathrm{G}_{2}$ and Sasakian geometry. According to the reductive decomposition of $V^{5,2}$, we provide an explicit description of all invariant $\mathrm{G}_{2}$ and Sasakian structures. In particular, we characterise the invariant $\mathrm{G}_{2}$-structures inducing a Sasakian metric, among which the well known nearly parallel $\mathrm{G}_{2}$-structure (Sasaki-Einstein) is included. As a consequence, we classify the invariant connections on homogeneous principal bundles over $V^{5,2}$ with gauge group ${\rm U}(1)$ and ${\rm SO}(3)$, satisfying either the $\mathrm{G}_{2}$ or the Sasakian instanton condition. Finally, we analyse the Yang Mills condition for those invariant connections.
@@ 1 Introduction
The Stiefel manifold $V^{n,k}$ of orthonormal $k$-frames in the $n$-dimensional Euclidean space (i.e. $1\leq k\leq n$) admits a transitive ${\rm SO}(n)$-action with isotropy subgroup isomorphic to ${\rm SO}(n-k)$, thus, it can be written homogeneously as $V^{n,k}={\rm SO}(n)/{\rm SO}(n-k)$ and consequently, its family of ${\rm SO}(n)$-invariant Riemannian metrics are described in terms of the isotropy representation of $V^{n,k}$. For instance, for $k=1$ or $k=n$, the isotropy representation is irreducible (indeed $V^{n,1}=S^{n}$ or $V^{n,n}={\rm SO}(n)$), in those cases, there is a unique ${\rm SO}(n)$-invariant metric, which is Einstein [wolf1968]. For $k=2$, the isotropy representation is reducible and the space of ${\rm SO}(n)$-invariant Riemannian metrics is described by a $4$-dimensional family, within there is a unique Einstein metric, which is inherited from the Grasmmanian $G_{2}(\mathbb{R}^{n})$ (e.g. [Arvanitoyeorgos1996, kerr1998new]). In addition, $V^{n,2}$ fits as an instance of the Boothby-Wang construction, as a total space of a $S^{1}$-principal bundle [boothby1958]*Theorem 3, moreover, it has a regular contact structure whose associated Reeb vector field generates a $1$-parameter subgroup of right translations of $V^{n,2}$.
A contact structure [^1] The reader unfamiliar with Sasakian geometry can find a complete study on this subject in [Boyer2008] on a $n$-dimensional manifold $M$ is a 1-form satisfying $\eta\wedge(d\eta)^{n}\neq 0$, in this case there exist a unique vector $\xi\in\mathscr{X}(M)$, called Reeb vector field, such that $\eta(\xi)=1$. The contact structure is called an almost contact structure if there exist an endomorphism $\Phi\in{\mathrm{End}}(TM)$ which satisfy $\Phi^{2}=-{\rm I}_{TM}+\eta\otimes\xi.$ A Riemannian metric $g$ is called compatible with the contact structure, if $g$ satisfies $g(\Phi X,\Phi Y)=g(X,Y)-\eta(X)\eta(Y),$ in this case, the contact structure is called an almost contact metric structure. In addition, if $\mathcal{S}=(\eta,\xi,g,\Phi)$ satisfies $2g(X,\Phi(Y))=d\eta(X,Y)$, the quadruple $\mathcal{S}$ is called a contact metric structure. Furthermore, if the Nijenhuis tensor $N_{\Phi}$ induced by $\Phi$ satisfies $N_{\Phi}=-d\eta\otimes\xi$, then $\mathcal{S}$ is called a Sasakian structure. In this paper, we consider the $7$–dimensional Stiefel manifold $V^{5,2}$ endowed with a one-parameter family of ${\rm SO}(5)$-invariant Sasakian structures $\mathcal{S}=(\eta,\xi,g,\Phi)$. For instance, homogeneous contact/Sasakian structures have been study in [andrada2009, correa2019] and [perrone1998homogeneous].
On the other hand, a differential $3$-form $\varphi$ on an oriented $7$–dimensional manifold $M$ is called a $\mathrm{G}_{2}$–structure if at each $p\in M$, there is a basis $e_{1},\dots,e_{7}$ of $T_{p}M$ such that
@@
where $e^{1},\dots,e^{7}$ denotes the dual basis of $T_{p}^{\ast}M$ and $e^{ijk}=e^{i}\wedge e^{j}\wedge e^{k}$. In particular, $\varphi$ determines a Riemannian metric $g_{\varphi}$, we study the family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$, such that the induced metric $g_{\varphi}$ is Sasakian. With respect to these structures, we study the theoretical Gauge instanton equations associated with the underlying Sasakian/$\mathrm{G}_{2}$-geometry of $V^{5,2}$. Namely, on a $G$–principal bundle $P\to M^{7}$ with $\sigma\in\Omega^{3}(M)$, the $1$-form connection $A\in\mathcal{A}(P)$ is called an $\lambda$-instanton, if it satisfies (e.g. [donaldson1998gauge, Tian2000])
@@
    A Self-dual contact Instantons (SDCI) [portilla2023] is a solutions of (1.2) with $\lambda=1$ and $\sigma:=\eta\wedge\omega$ naturally defined by a contact structure of $(M,\eta)$, where $\omega:=\frac{1}{2}d\eta$ is called the fundamental $2$-form..
@@
    A $\mathrm{G}_{2}$–instantons [corrigan1983first, donaldson1998gauge] is a solutions of (1.2) with $\lambda=-1$ and $\sigma=\varphi$ a coclosed (i.e. $d\ast\varphi=0$) $\mathrm{G}_{2}$–structure on $M$.
Furthermore, for $E\to(M^{7},\mathcal{S})$ a Sasakian holomorphic bundle [biswas2010vector], the usual notion of Hermitian Yang-Mills (HYM) connection on a Kähler manifolds is extended to odd dimension in the natural fashion to the so-called transverse Hermitian Yang-Mills connections (tHYM) (see (3.1)). The concept of Chern connections also extend naturally to a Sasakian version, as a connection mutually compatible with the transverse holomorphic structure and the Hermitian metric. A noteworthy result is that if $E\to M$ is a holomorphic Sasakian bundle on a contact Calabi-Yau manifold $M$, along compatible connections the notion of SDCI, $\mathrm{G}_{2}$-instantons and tHYM coincide. cf. [portilla2023, Teorem 1.1] and [calvo2020gauge, Lemma  21].
In [donaldson1998gauge, Donaldson2011], Donaldson, Segal and Thomas suggested that it may be possible to use $\mathrm{G}_{2}$-instantons to construct an enumerative invariant of $\mathrm{G}_{2}$-manifolds. For instance, when $(M^{7},\varphi)$ has a torsion-free $\mathrm{G}_{2}$-structure $\varphi$ (i.e. $d\varphi=0$ and $d\ast\varphi=0$ [fernandez1982riemannian]), solutions to the $\mathrm{G}_{2}$-instanton equation have been studied in [sa2015g2, walpuski2015g_2, clarke2014instantons, lotay20182]. However, the construction of metrics with holonomy $\mathrm{G}_{2}$ has proven to be a challenging task, thus, the torsion-free condition has been relaxed to the class of coclosed $\mathrm{G}_{2}$-structures [ball2019gauge, clarke2023], which has a notably application in theoretical physics to the Killing spinor equations in supergravity (cf. [Ivanov2005, Ivanov10, delaOssaG21, Lotay2022, Fernandez2011]). Analogously, Wang has proposed studying the compactification of the moduli space of contact instantons as a tool to construct a Donaldson-type invariant for contact $7$-manifolds [wang2012higher].
@@
In particular, it defines a family of (non-purely) coclosed $\mathrm{G}_{2}$-structures, within the nearly parallel $\mathrm{G}_{2}$-structures arises as an $S^{1}$-family for $x=a$, $y=b$ and $a^{2}+b^{2}=z^{2}=\frac{27}{512}$ (cf. Theorem 2.1).
@@
In Section 3, we revisit the interaction between SDCI, $\mathrm{G}_{2}$-instantons and tHYM connections addressed in [portilla2023, Proposition 3.11] and [calvo2020gauge, Lemma 21] against the backdrop of contact Calabi-Yau manifolds. For a bundle $E\to V^{5,2}$, in Lemma 3.1 we prove that if the connection $A$ is SDCI then it is a $\mathrm{G}_{2}$-instanton. Conversely, for a holomorphic Sasakian bundle $E\to V^{5,2}$, if the Chern connection is a $\mathrm{G}_{2}$-instanton, then it is a SCDI. Finally, we prove that HYM connections on a holomorphic bundle $E\to G_{2}(\mathbb{R}^{5})$ on the Grassmanian, pulling back to $\mathrm{G}_{2}$-instantons on $\pi^{\ast}E\to V^{5,2}$ (cf. Corollary 3.4). We recall here that the key argument follow just in presence of a SU$(3)$-structure.
Finally, in Section 4 we provide examples of homogeneous SCDI and $\mathrm{G}_{2}$-instantons over $V^{5,2}$, with gauge group ${\rm SO}(3)$ and ${\rm U}(1)$. In particular, we prove that the ${\rm SO}(3)$-invariant $\mathrm{G}_{2}$-instanton (4.8) is a Yang-Mills connection on $(V^{5,2},\varphi)$, for any ${\rm SO}(5)$-invariant coclosed $\mathrm{G}_{2}$-structure (cf. Theorem 4.1), moreover, it describes a local minimizer of the Yang-Mills functional (cf. Corollary 4.11).
Outline: In Section 2, we provide some preliminaries about $V^{5,2}$ as a homogeneous space, with the aim of characterising all the ${\rm SO}(5)$-invariant coclosed $\mathrm{G}_{2}$-structures and the associated Sasakian metrics. In Section 3, we explore the interaction between SDCI and $\mathrm{G}_{2}$–instantons on $V^{5,2}$, in particular, Lemmata 3.1, 3.3 and Corollary 3.4 indicate conditions such that contact instantons and $\mathrm{G}_{2}$-instantons coincides. Finally, in Section 4 we consider homogeneous principal bundles on $V^{5,2}$ and we construct examples of invariant instantons on ${\rm SO}(3)$, $U(1)$–homogeneous bundles. In particular, in Section 4.1 we characterise all invariant ${\rm SO}(3)$–connection and we show that these satisfy the $\mathrm{G}_{2}$ and SDCI condition for appropriate parameters and we analyse its stability as Yang-Mills connections.
Acknowledgments and funding: AM was funded by the São Paulo Research Foundation (Fapesp) [2021/08026-5] and [2023/13780-6]. LP was funded by Centre Henri Lebesgue (CHL) under the grant ANR-11-LABX-0020-01, he also would like to thanks to LMBA at Université de Bretagne Occidentale UBO for the hospitality.
@@ 2 Homogeneous $\mathrm{G}_{2}$ and contact metric structures on $V^{5,2}$
In this section, we provide an explicit description of a $\mathrm{Ad}({\rm SO}(3))$-invariant $\mathrm{G}_{2}$-structure and the induced inner product on $V^{5,2}$. As well as, we establish the notation that we will use throughout this work.
@@ 2.1 $\mathrm{G}_{2}$-structures on $V^{5,2}$
In order to describe the space of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structure on $V^{5,2}={\rm SO}(5)/{\rm SO}(3)$ [^2] This subject is brought up in [friedrich1997nearly]., we fix a reductive decomposition of $V^{5,2}$ as a homogeneous space. Consider the Lie algebra of ${\rm SO}(5)$
@@
and denote by $E_{ij}$, the $5\times 5$ matrix with $1$ at the $(ij)$-entry and 00 elsewhere, and fix a basis for $\mathfrak{so}(5)$:
$\begin{split}e_{1}=&E_{12}-E_{21},\quad e_{4}=E_{15}-E_{51},\quad e_{7}=E_{25}-E_{52},\quad e_{10}=E_{54}-E_{45},\\ e_{2}=&E_{13}-E_{31},\quad e_{5}=E_{23}-E_{32},\quad e_{8}=E_{34}-E_{43},\quad\\ e_{3}=&E_{14}-E_{41},\quad e_{6}=E_{24}-E_{42},\quad e_{9}=E_{35}-E_{53}.\end{split}$ (2.2)
We consider the Lie algebra $\mathfrak{so}(3)$ embedded in $\mathfrak{so}(5)$ and it is spanned by $\{e_{8},e_{9},e_{10}\}$. The Lie bracket of $\mathfrak{so}(5)$ is given by the commutator describe in Table 1.
[,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] $e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $e_{6}$ $e_{7}$ $e_{8}$ $e_{9}$ $e_{10}$
$e_{1}$ 00 $-e_{5}$ $-e_{6}$ $-e_{7}$ $e_{2}$ $e_{3}$ $e_{4}$ 00 00 00
$e_{2}$ $e_{5}$ 00 $-e_{8}$ $-e_{9}$ $-e_{1}$ 00 00 $e_{3}$ $e_{4}$ 00
$e_{3}$ $e_{6}$ $e_{8}$ 00 $-e_{10}$ 00 $-e_{1}$ 00 $-e_{2}$ 00 $e_{4}$
$e_{4}$ $e_{7}$ $e_{9}$ $e_{10}$ 00 00 00 $-e_{1}$ 00 $-e_{2}$ $-e_{3}$
$e_{5}$ $-e_{2}$ $e_{1}$ 00 00 00 $-e_{8}$ $-e_{9}$ $e_{6}$ $e_{7}$ 00
$e_{6}$ $-e_{3}$ 00 $e_{1}$ 00 $e_{8}$ 00 $-e_{10}$ $-e_{5}$ 00 $e_{7}$
$e_{7}$ $-e_{4}$ 00 00 $e_{1}$ $e_{9}$ $e_{10}$ 00 00 $-e_{5}$ $-e_{6}$
$e_{8}$ 00 $-e_{3}$ $e_{2}$ 00 $-e_{6}$ $e_{5}$ 00 00 $-e_{10}$ $e_{9}$
$e_{9}$ 00 $-e_{4}$ 00 $e_{2}$ $-e_{7}$ 00 $e_{5}$ $e_{10}$ 00 $-e_{8}$
$e_{10}$ 00 00 $-e_{4}$ $e_{3}$ 00 $-e_{7}$ $e_{6}$ $-e_{9}$ $e_{8}$ 00
@@
$\mathrm{ad}(e_{10})_{\mathfrak{m}}$ $=\mathrm{diag}(0,A_{3}A_{3}),$
where $A_{1}=\begin{pmatrix}0&-1&0\\ 1&0&0\\ 0&0&0\end{pmatrix}$,  $A_{2}=\begin{pmatrix}0&0&-1\\ 0&0&0\\ 1&0&0\end{pmatrix}$ and $A_{3}=\begin{pmatrix}0&0&0\\ 0&0&-1\\ 0&1&0\end{pmatrix}$ are the canonical basis of $\mathfrak{so}(3)$. Denote by ${\mathfrak{m}}_{1}=\operatorname{{\rm Span}}\{e_{1}\},$ ${\mathfrak{m}}_{2}:=\operatorname{{\rm Span}}\{e_{2},e_{3},e_{4}\}$ and ${\mathfrak{m}}_{3}:=\operatorname{{\rm Span}}\{e_{5},e_{6},e_{7}\}.$ By using the Lie bracket in Table 1 is straightforward to verify that
@@
$\begin{array}[]{rl}\omega_{0}&=e^{25}+e^{36}+e^{47},\\[3.0pt] \mathop{\mathrm{Im}}\Omega_{0}&=e^{237}-e^{246}+e^{345}-e^{567},\\[3.0pt] \mathop{\mathrm{Re}}\Omega_{0}&=e^{234}-e^{267}+e^{357}-e^{456}.\end{array}$ (2.6)
Notice that $\omega_{0}\wedge\mathop{\mathrm{Re}}\Omega_{0}=0$ and $2\omega_{0}^{3}=3\mathop{\mathrm{Re}}\Omega_{0}\wedge\mathop{\mathrm{Im}}\Omega_{0}\neq 0$, i.e. $(\omega_{0},\mathop{\mathrm{Re}}\Omega_{0})$ is a ${\rm SU}(3)$-structure on ${\mathfrak{p}}$ (with induced orientation $-e^{234567}$).
@@
The forms in (2.6) are $\mathrm{ad}(\mathfrak{so}(2)\oplus\mathfrak{so}(3))$-invariant, therefore, it induces a ${\rm SO}(5)$-invariant ${\rm SU}(3)$-structure $(\omega,\mathop{\mathrm{Re}}\Omega)$ on $G_{2}(\mathbb{R}^{5})$. Furthermore, we have that $d\omega=0,$ $d(\mathop{\mathrm{Re}}\Omega)=3e^{1}\wedge\mathop{\mathrm{Im}}\Omega$ and $d(\mathop{\mathrm{Im}}\Omega)=-3e^{1}\wedge\mathop{\mathrm{Re}}\Omega.$

### v2

@@ Abstract
We study homogeneous instantons on the seven–dimensional Stiefel manifold $V^{5,2}$ in the context of $\mathrm{G}_{2}$ and Sasakian geometry. According to the reductive decomposition of $V^{5,2}$, we provide an explicit description of all invariant $\mathrm{G}_{2}$ and Sasakian structures. In particular, we characterise the invariant $\mathrm{G}_{2}$–structures inducing a Sasakian metric, among which the well–known nearly parallel $\mathrm{G}_{2}$–structure (Sasaki–Einstein) is included. As a consequence, we classify the invariant connections on homogeneous principal bundles over $V^{5,2}$ with gauge group ${\rm U}(1)$ and ${\rm SO}(3)$, satisfying either the $\mathrm{G}_{2}$ or the Sasakian instanton condition. In addition, we study infinitesimal deformations of $\mathrm{G}_{2}$–instantons on coclosed $\mathrm{G}_{2}$–manifolds using a spinorial approach. By means of a Weitzenböck–type formula with torsion, we obtain curvature obstructions to the existence of non–trivial infinitesimal deformations and prove rigidity results for certain homogeneous $\mathrm{G}_{2}$–instantons.
@@ 1 Introduction
The Stiefel manifold $V^{n,k}$ of orthonormal $k$-frames in the $n$-dimensional Euclidean space (i.e. $1\leq k\leq n$) admits a transitive ${\rm SO}(n)$-action with isotropy subgroup isomorphic to ${\rm SO}(n-k)$, thus, it can be written homogeneously as $V^{n,k}={\rm SO}(n)/{\rm SO}(n-k)$ and consequently, its family of ${\rm SO}(n)$-invariant Riemannian metrics is described in terms of the isotropy representation of $V^{n,k}$. For instance, for $k=1$ or $k=n$, the isotropy representation is irreducible (indeed $V^{n,1}=S^{n}$ or $V^{n,n}={\rm SO}(n)$), in those cases, there is a unique ${\rm SO}(n)$-invariant metric, which is Einstein [wolf1968]. For $k=2$, the isotropy representation is reducible and the space of ${\rm SO}(n)$-invariant Riemannian metrics is described by a $4$-dimensional family, within which there exists a unique Einstein metric, which is inherited from the Grassmannian $G_{2}(\mathbb{R}^{n})$ (e.g. [Arvanitoyeorgos1996, kerr1998new]). In addition, $V^{n,2}$ fits as an instance of the Boothby-Wang construction, as the total space of a $S^{1}$-principal bundle [boothby1958]*Theorem 3, moreover, it has a regular contact structure whose associated Reeb vector field generates a one-parameter subgroup of right translations of $V^{n,2}$.
A contact structure [^1] The reader unfamiliar with Sasakian geometry can find a complete study on this subject in [Boyer2008] on a $(2n+1)$-dimensional manifold $M$ is a 1-form satisfying $\eta\wedge(d\eta)^{n}\neq 0$, in this case there exists a unique vector $\xi\in\mathscr{X}(M)$, called Reeb vector field, such that $\eta(\xi)=1$. The contact structure is called an almost contact structure if there exists an endomorphism $\Phi\in{\mathrm{End}}(TM)$ which satisfy $\Phi^{2}=-{\rm I}_{TM}+\eta\otimes\xi.$ A Riemannian metric $g$ is called compatible with the contact structure, if $g$ satisfies $g(\Phi X,\Phi Y)=g(X,Y)-\eta(X)\eta(Y),$ in this case, the contact structure is called an almost contact metric structure. In addition, if $\mathcal{S}=(\eta,\xi,g,\Phi)$ satisfies $2g(X,\Phi(Y))=d\eta(X,Y)$, the quadruple $\mathcal{S}$ is called a contact metric structure. Furthermore, if the Nijenhuis tensor $N_{\Phi}$ induced by $\Phi$ satisfies $N_{\Phi}=-d\eta\otimes\xi$, then $\mathcal{S}$ is called a Sasakian structure. In this paper, we consider the 77–dimensional Stiefel manifold $V^{5,2}$ endowed with a one-parameter family of ${\rm SO}(5)$-invariant Sasakian structures $\mathcal{S}=(\eta,\xi,g,\Phi)$. For instance, homogeneous contact/Sasakian structures have been studied in [andrada2009, correa2019] and [perrone1998homogeneous].
On the other hand, a differential $3$-form $\varphi$ on an oriented 77–dimensional manifold $M$ is called a $\mathrm{G}_{2}$–structure if at each $p\in M$, there is a basis $e_{1},\dots,e_{7}$ of $T_{p}M$ such that
@@
where $e^{1},\dots,e^{7}$ denotes the dual basis of $T_{p}^{\ast}M$ and $e^{ijk}=e^{i}\wedge e^{j}\wedge e^{k}$. In particular, $\varphi$ determines a Riemannian metric $g_{\varphi}$, we study the family of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}$, such that the induced metric $g_{\varphi}$ is Sasakian. With respect to these structures, we study the theoretical gauge instanton equations associated with the underlying Sasakian and $\mathrm{G}_{2}$-geometry of $V^{5,2}$. Namely, on a GG–principal bundle $P\to M^{7}$ with $\sigma\in\Omega^{3}(M)$, the $1$-form connection $A\in\mathcal{A}(P)$ is called an $\lambda$-instanton, if it satisfies (e.g. [donaldson1998gauge, Tian2000])
@@
    A Self-dual contact instantons (SDCI) [portilla2023] is a solution of (1.2) with $\lambda=1$ and $\sigma:=\eta\wedge\omega$ naturally defined by a contact structure of $(M,\eta)$, where $\omega:=\frac{1}{2}d\eta$ is called the fundamental $2$-form.
@@
    A $\mathrm{G}_{2}$–instantons [corrigan1983first, donaldson1998gauge] is a solution of (1.2) with $\lambda=-1$ and $\sigma=\varphi$ a coclosed (i.e. $d\ast\varphi=0$) $\mathrm{G}_{2}$–structure on $M$. Equivalently, a connection $A$ is a $\mathfrak{g}$}–instanton if and only if
    $F_{A}\wedge\ast\varphi=0$ (1.3)
Furthermore, for $E\to(M^{7},\mathcal{S})$ a Sasakian holomorphic bundle [biswas2010vector], the usual notion of Hermitian Yang-Mills (HYM) connection on a Kähler manifolds is extended to odd dimension in the natural fashion to the so-called transverse Hermitian Yang-Mills connections (tHYM) (see (3.1)). The concept of Chern connections also extends naturally to a Sasakian version, as a connection mutually compatible with the transverse holomorphic structure and the Hermitian metric. A noteworthy result is that if $E\to M$ is a holomorphic Sasakian bundle on a contact Calabi-Yau manifold $M$, along compatible connections the notion of SDCI, $\mathrm{G}_{2}$-instantons and tHYM coincide. cf. [portilla2023, Theorem 1.1] and [calvo2020gauge, Lemma  21].
In [donaldson1998gauge, Donaldson2011], Donaldson, Segal and Thomas suggested that it may be possible to use $\mathrm{G}_{2}$-instantons to construct an enumerative invariant of $\mathrm{G}_{2}$-manifolds. For instance, when $(M^{7},\varphi)$ has a torsion-free $\mathrm{G}_{2}$-structure $\varphi$ (i.e. $d\varphi=0$ and $d\ast\varphi=0$ [fernandez1982riemannian]), solutions to the $\mathrm{G}_{2}$-instanton equation have been studied in [sa2015g2, walpuski2015g_2, clarke2014instantons, lotay20182, singhal2022]. However, the construction of metrics with holonomy $\mathrm{G}_{2}$ has proven to be a challenging task, thus, the torsion-free condition has been relaxed to the class of coclosed $\mathrm{G}_{2}$-structures [ball2019gauge, clarke2023], which has a notable application in theoretical physics to the Killing spinor equations in supergravity (cf. [Ivanov2005, Ivanov10, delaOssaG21, Lotay2022, Fernandez2011]). Analogously, Wang has proposed studying the compactification of the moduli space of contact instantons as a tool to construct a Donaldson-type invariant for contact $7$-manifolds [wang2012higher].
@@
In particular, it defines a family of (non-purely) coclosed $\mathrm{G}_{2}$-structures, within the nearly parallel $\mathrm{G}_{2}$-structures arises as an $S^{1}$–family for $x=a$, $y=b$ and $a^{2}+b^{2}=z^{2}=\frac{27}{512}$ (cf. Theorem 2.1).
@@
In Section 3, we revisit the interaction between SDCI, $\mathrm{G}_{2}$-instantons and tHYM connections addressed in [portilla2023, Proposition 3.11] and [calvo2020gauge, Lemma 21] against the backdrop of contact Calabi-Yau manifolds. For a bundle $E\to V^{5,2}$, in Lemma 3.1 we prove that if the connection $A$ is SDCI then it is a $\mathrm{G}_{2}$-instanton. Conversely, for a holomorphic Sasakian bundle $E\to V^{5,2}$, if the Chern connection is a $\mathrm{G}_{2}$-instanton, then it is a SCDI. Finally, we prove that HYM connections on a holomorphic bundle $E\to G_{2}(\mathbb{R}^{5})$ on the Grassmannian, pulling back to $\mathrm{G}_{2}$-instantons on $\pi^{\ast}E\to V^{5,2}$ (cf. Corollary 3.4). We point out that the key argument can be carried out under the sole assumption of an $\mathrm{SU}(3)$-structure, thereby improving the results mentioned above.
In Section 4, we provide examples of homogeneous SDCI and $\mathrm{G}_{2}$-instantons over $V^{5,2}$, with gauge group ${\rm SO}(3)$ and ${\rm U}(1)$. In particular, we prove that the ${\rm SO}(3)$-invariant $\mathrm{G}_{2}$-instanton (4.9) is a Yang-Mills connection on $(V^{5,2},\varphi)$, for any ${\rm SO}(5)$-invariant coclosed $\mathrm{G}_{2}$-structure (cf. Theorem 4.1).
Finally, in Section 5 we extend the spinorial framework proposed by R. Singhal [singhal2022] to coclosed $\mathrm{G}_{2}$–structures, and as a concrete application in Corollary 5.7, we prove rigidity for a family of homogeneous $\mathrm{G}_{2}$–instantons on the Stiefel manifold $V^{5,2}$ for a non-empty range of the metric parameter $y_{2}$.
Outline: In Section 2, we provide some preliminaries about $V^{5,2}$ as a homogeneous space, with the aim of characterising all the ${\rm SO}(5)$-invariant coclosed $\mathrm{G}_{2}$-structures and the associated Sasakian metrics. In Section 3, we explore the interaction between SDCI and $\mathrm{G}_{2}$–instantons on $V^{5,2}$, in particular, Lemmata 3.1, 3.3 and Corollary 3.4 indicate conditions such that contact instantons and $\mathrm{G}_{2}$-instantons coincide. Finally, in Section 4 we consider homogeneous principal bundles on $V^{5,2}$ and we construct examples of invariant instantons on ${\rm SO}(3)$, U(1)U(1)–homogeneous bundles. In particular, in Section 4.1 we characterise all invariant ${\rm SO}(3)$–connection and we show that these satisfy the $\mathrm{G}_{2}$ and SDCI condition for appropriate parameters. Finally, in Section 5 we study the deformation theory for a particular example of a $\mathrm{G}_{2}$-instanton, specifically in Corollary 5.7 we prove that it is rigid for a non-empty interval of the metric parameter $y_{2}$.
Acknowledgments and funding: AM was funded by the São Paulo Research Foundation [2021/08026-5] and [2023/13780-6] (Fapesp). LP was funded by Centre Henri Lebesgue (CHL) under the grant ANR-11-LABX-0020-01. He also thanks the LMBA at Université de Bretagne Occidentale (UBO) for its hospitality
@@ 2 Homogeneous $\mathrm{G}_{2}$ and contact metric structures on $V^{5,2}$
In this section, we provide an explicit description of a $\mathrm{Ad}({\rm SO}(3))$-invariant $\mathrm{G}_{2}$-structure and the induced inner product on $V^{5,2}$. We also establish the notation that we will use throughout this work.
@@ 2.1 $\mathrm{G}_{2}$-structures on $V^{5,2}$
In order to describe the space of ${\rm SO}(5)$-invariant $\mathrm{G}_{2}$-structures on $V^{5,2}={\rm SO}(5)/{\rm SO}(3)$ [^2] This subject is discussed in depth in [friedrich1997nearly]., we fix a reductive decomposition of $V^{5,2}$ as a homogeneous space. Consider the Lie algebra of ${\rm SO}(5)$
@@
Let $E_{ij}$ denote the elementary $5\times 5$ matrix with a $1$ in the $(i,j)$-entry and zeros elsewhere. We fix the following basis of the Lie algebra $\mathfrak{so}(5)$:
$\begin{split}e_{1}=&E_{12}-E_{21},\quad e_{4}=E_{15}-E_{51},\quad e_{7}=E_{25}-E_{52},\quad e_{10}=E_{54}-E_{45},\\[3.00003pt] e_{2}=&E_{13}-E_{31},\quad e_{5}=E_{23}-E_{32},\quad e_{8}=E_{34}-E_{43},\quad\\[3.00003pt] e_{3}=&E_{14}-E_{41},\quad e_{6}=E_{24}-E_{42},\quad e_{9}=E_{35}-E_{53}.\end{split}$ (2.2)
We consider the Lie algebra $\mathfrak{so}(3)$ embedded in $\mathfrak{so}(5)$ which is spanned by $\{e_{8},e_{9},e_{10}\}$. The Lie bracket of $\mathfrak{so}(5)$ is given by the commutator described in Table 1.
$\cdot$$\cdot$] $e_{1}$ $e_{2}$ $e_{3}$ $e_{4}$ $e_{5}$ $e_{6}$ $e_{7}$ $e_{8}$ $e_{9}$ $e_{10}$
$e_{1}$ 0 $e_{5}$ $e_{6}$ $e_{7}$ $e_{2}$ $e_{3}$ $e_{4}$ 0 0 0
$e_{2}$ $e_{5}$ 0 $e_{8}$ $e_{9}$ $e_{1}$ 0 0 $e_{3}$ $e_{4}$ 0
$e_{3}$ $e_{6}$ $e_{8}$ 0 $e_{10}$ 0 $e_{1}$ 0 $e_{2}$ 0 $e_{4}$
$e_{4}$ $e_{7}$ $e_{9}$ $e_{10}$ 0 0 0 $e_{1}$ 0 $e_{2}$ $e_{3}$
$e_{5}$ $e_{2}$ $e_{1}$ 0 0 0 $e_{8}$ $e_{9}$ $e_{6}$ $e_{7}$ 0
$e_{6}$ $e_{3}$ 0 $e_{1}$ 0 $e_{8}$ 0 $e_{10}$ $e_{5}$ 0 $e_{7}$
$e_{7}$ $e_{4}$ 0 0 $e_{1}$ $e_{9}$ $e_{10}$ 0 0 $e_{5}$ $e_{6}$
$e_{8}$ 0 $e_{3}$ $e_{2}$ 0 $e_{6}$ $e_{5}$ 0 0 $e_{10}$ $e_{9}$
$e_{9}$ 0 $e_{4}$ 0 $e_{2}$ $e_{7}$ 0 $e_{5}$ $e_{10}$ 0 $e_{8}$
$e_{10}$ 0 0 $e_{4}$ $e_{3}$ 0 $e_{7}$ $e_{6}$ $e_{9}$ $e_{8}$ 0
@@
$\mathrm{ad}(e_{10})_{\mathfrak{m}}$ $=\mathrm{diag}(0,A_{3},A_{3}),$
where $A_{1}=\begin{pmatrix}0&-1&0\\ 1&0&0\\ 0&0&0\end{pmatrix}$,   $A_{2}=\begin{pmatrix}0&0&-1\\ 0&0&0\\ 1&0&0\end{pmatrix}$ and $A_{3}=\begin{pmatrix}0&0&0\\ 0&0&-1\\ 0&1&0\end{pmatrix}$ are the canonical basis of $\mathfrak{so}(3)$. Denote by ${\mathfrak{m}}_{1}=\operatorname{{\rm Span}}\{e_{1}\},$ ${\mathfrak{m}}_{2}:=\operatorname{{\rm Span}}\{e_{2},e_{3},e_{4}\}$ and ${\mathfrak{m}}_{3}:=\operatorname{{\rm Span}}\{e_{5},e_{6},e_{7}\}.$ By using the Lie bracket in Table 1, it is straightforward to verify that
@@
$\begin{array}[]{rl}\omega_{0}&=e^{25}+e^{36}+e^{47},\\[4.0pt] \mathop{\mathrm{Im}}\Omega_{0}&=e^{237}-e^{246}+e^{345}-e^{567},\\[4.0pt] \mathop{\mathrm{Re}}\Omega_{0}&=e^{234}-e^{267}+e^{357}-e^{456}.\end{array}$ (2.6)
Notice that $\omega_{0}\wedge\mathop{\mathrm{Re}}\Omega_{0}=0$ and $2\omega_{0}^{3}=3\mathop{\mathrm{Re}}\Omega_{0}\wedge\mathop{\mathrm{Im}}\Omega_{0}\neq 0$, i.e. $(\omega_{0},\mathop{\mathrm{Re}}\Omega_{0})$ is a ${\rm SU}(3)$-structure on $\mathfrak{p}$} (with induced orientation $e^{234567}$).
@@
The forms in (2.6) are $\mathrm{ad}(\mathfrak{so}(2)\oplus\mathfrak{so}(3))$-invariant, therefore, they induce a ${\rm SO}(5)$-invariant ${\rm SU}(3)$-structure $(\omega,\mathop{\mathrm{Re}}\Omega)$ on $G_{2}(\mathbb{R}^{5})$. Furthermore, we have that $d\omega=0,$ $d(\mathop{\mathrm{Re}}\Omega)=3e^{1}\wedge\mathop{\mathrm{Im}}\Omega$ and $d(\mathop{\mathrm{Im}}\Omega)=-3e^{1}\wedge\mathop{\mathrm{Re}}\Omega.$